# Robust Multi-Agent Safety via Tube-Based Tightened Exponential Barrier Functions

### v1

@@ 1 Introduction
To operate successfully in applications like vehicle convoys and aerial swarms, multi-agent systems (MAS) must satisfy strict coordination and safety demands (Barnes et al., 2007). This challenge is amplified for systems with high relative degrees, where the chain of integrators between the control input and the safety output introduces significant response latency, heightening the risk of collisions in safety-critical settings (Nguyen and Sreenath, 2016; Garg et al., 2024). Model Predictive Control (MPC) offers a powerful paradigm for constrained control by optimizing over future predictions (Mayne et al., 2000). In parallel, higher-order and exponential control barrier functions (HOCBFs and eCBFs) have emerged as effective tools for enforcing safety in systems with complex dynamics (Xiao and Belta, 2022, 2019; Tan et al., 2022; Nguyen and Sreenath, 2016; Wang et al., 2021). However, enforcing robust safety for multi-agent systems in the presence of model uncertainty and external disturbances remains a significant challenge, especially when these advanced CBF techniques are applied in a distributed context.
@@
This work introduces a robust safety framework for leader-follower formation control in continuous-time, nonlinear multi-agent systems subject to bounded matched disturbances. The central contribution is a method for formally coupling robust error feedback with nominal trajectory planning to enforce safety specifications. Our approach addresses a critical deficiency in the literature: while methods combining distributed MPC with HOCBFs exist (Wang et al., 2025), their safety analysis is often confined to nominal, disturbance-free models and lacks formal mechanisms to ensure forward invariance for systems under uncertainty. Similarly, other distributed control schemes for multi-agent systems (Tan and Dimarogonas, 2022; Mestres et al., 2024) are often reactive and lack the foresight of a planning horizon, or, if predictive (Jiang and Guo, 2024), have not been formulated to provide formal robustness guarantees for systems with high relative-degree dynamics. By integrating robust tube constructions (Kolathaya and Ames, 2019) with techniques for high-order constraints (Xiao and Belta, 2019; Ames et al., 2017), the presented framework provides a unified solution that simultaneously offers predictive planning, robust safety, and applicability to high-relative-degree systems.
We present a control synthesis framework that seamlessly integrates robust feedback with nominal planning. Local ancillary feedback controllers first confine the deviations of the true agent states from their nominal counterparts within pre-computed robust positively invariant (RPI) tubes. The key insight is that the geometries of these tubes provide sufficient information to systematically tighten the exponential control barrier function (eCBF) constraints imposed on the nominal planner of each agent. This integrated synthesis ensures that any set of nominal trajectories satisfying the tightened constraints corresponds to a provably safe trajectory for the true, disturbed system.
To leverage this synthesis for high-performance planning, we implement the planner using a distributed Model Predictive Control (DMPC) framework. Unlike reactive controllers that rely only on the current state, the predictive horizon of DMPC allows each agent to proactively shape its trajectory to avoid future hazards and inevitable collisions, a known limitation of single-timestep safety methods. The tightened eCBF inequalities are directly incorporated into each agent’s local optimization problem, resulting in a computationally tractable controller that combines the foresight of MPC with formal, end-to-end guarantees of robust safety and recursive feasibility.
The remainder of the paper is structured as follows: Section 2 presents the multi-agent dynamics, derives the nominal control model, constructs per-agent RPI tubes around the nominal trajectories, and defines support functions to tighten the nominal eCBF constraints. Section 3 provides the tube-MPC with tightened eCBF inequality framework over fixed communication topology. Section 4 demonstrates the proposed framework via a numerical example with fixed communication topology. Section 5 summarizes contributions and potential future work.
@@ 2.1 State Based Ancillary Feedback:
We assume a Lipschitz bound on the nonlinearity around the nominal trajectory
$\|\Delta f^{\,i}(\bar{x}^{\,i},z^{\,i},t)\|\ =\big\|f^{\,i}(\bar{x}^{\,i}\!+z^{\,i},t)-f^{\,i}(\bar{x}^{\,i},t)\big\|\leq\ L_{f^{i}}\,\|z^{\,i}\|,$
with a known constant $L_{f^{i}}\geq 0$.
@@ 2.2 Robust Positive Invariant (RPI) Tube Construction:
$\rho_{i}\geq\frac{2\bar{w}_{i}\lambda_{\max}(P_{i})}{\lambda_{\min}(Q_{i})-2L_{f^{i}}\lambda_{\max}(P_{i})}.$
$\square$
@@
Since $\|w^{i}(t)\|\leq\bar{w}_{i}$, a necessary and sufficient condition is
@@ 2.3 Tightened Safety and Support Functions
where $F_{x}(x^{i})=\begin{bmatrix}{x}_{2}^{i^{\top}},\cdots,f^{i}(x^{i},t)^{\top}\end{bmatrix}^{\top}$, $F_{u}=\begin{bmatrix}0,0,\cdots,I\end{bmatrix}^{\top}\!$, $F_{w}=\begin{bmatrix}0,0,\cdots,I\end{bmatrix}^{\top}$. Accordingly, the nominal control–affine dynamics of agent $i$ from (2) is written
@@
$\mathcal{C}:=\bigg\{\left(x^{1},\cdots,x^{i},\cdots,x^{N}\right)\in\mathbb{R}^{N\cdot n\cdot d}:\\ \hfill h_{ij}(x^{i},x^{j})\geq 0,\hskip 20.0pt\forall(i,j)\in\mathcal{V}^{2},\hskip 45.0pt\\ h_{iO}(x^{i})\geq 0,\hskip 20.0pt\forall(i,O)\in\mathcal{V}\times\mathcal{O}_{i}\bigg\},$ (20)
@@
Assumption 2.1
We assume that $h_{ij}({x}^{i},{x}^{j})$, $h_{ij}(\bar{x}^{i},\bar{x}^{j})$, $h_{iO}({x}^{i})$ and $h_{iO}(\bar{x}^{i})$ are convex and continuously differentiable in their arguments. $\square$
@@
Given the decomposition $x^{\,i}=\bar{x}^{\,i}+z^{\,i}$, $x^{\,j}=\bar{x}^{\,j}+z^{\,j}$ with $z^{\,i}\in\mathcal{Z}_{i}$, $z^{\,j}\in\mathcal{Z}_{j}$, and $z^{ij}=z^{i}-z^{j}$, we define the inter-agent tightened safety function as the worst-case value (infimum) considering all possible errors in relative tube $\mathcal{Z}_{ij}=\mathcal{Z}_{i}\oplus(-\mathcal{Z}_{j})$ in the form of
@@
Lemma 2.3 (Support Function of Tubes)
The support function of the safety tube $\mathcal{Z}_{i}:=\{\,z^{\,i}:~(z^{\,i})^{\top}P_{i}z^{\,i}\leq\rho_{i}^{\,2}\,\}$, with $\rho_{i}$ and $P_{i}\succ 0$ satisfying the Lyapunov equation (9) as in Lemma 2.1, is
@@ 3.1 Proposed Tightened eCBF Functions
We propose the following tightened exponential control barrier functions (tight-eCBF) using the standard eCBF (3)–(3), to impose on the nominal trajectories of agents while ensuring safety of the actual trajectories which are subject to unknown disturbances $w^{i}$ and corresponding errors $z^{i}$.
@@
where we use notations $\delta_{ij}(\bar{x}^{i},\bar{x}^{j}):=\sigma_{\mathcal{Z}_{i}}(g_{i})+\sigma_{\mathcal{Z}_{j}}(g_{j})\equiv\rho_{i}\left\|\nabla_{\bar{x}^{i}}h_{ij}(\bar{x}^{i},\bar{x}^{j})\right\|_{P_{i}^{-1}}+\rho_{j}\left\|\nabla_{\bar{x}^{j}}h_{ij}(\bar{x}^{i},\bar{x}^{j})\right\|_{P_{j}^{-1}}$, together with $\delta_{iO}(\bar{x}^{i}):=\sigma_{\mathcal{Z}_{i}}(g_{iO})\equiv{\rho_{i}}\left\|\nabla_{\bar{x}^{i}}h_{iO}(\bar{x}^{i},\bar{x}^{j})\right\|_{P_{i}^{-1}}$.
@@
where $G_{i}^{\max}$ and $G_{j}^{\max}$ are finite bounds on the gradients of $h_{ij}$ with respect to $\bar{x}^{i}$ and $\bar{x}^{i}$, i.e.,
$|\nabla_{\bar{x}^{i}}h_{ij}(\bar{x}^{i},\bar{x}^{j})|\leq G_{i}^{\max},\qquad|\nabla_{\bar{x}^{j}}h_{ij}(\bar{x}^{i},\bar{x}^{j})|\leq G_{j}^{\max}.$
Hence
@@ 3.2 Distributed Tube-MPC eCBF Formation Framework
In order to reduce the amount of computations at the agent level, and to establish distributed control policies, we consider a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E},A)$ with the node set $\mathcal{V}=\{1,\dots,N\}$, edge set $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$, and weighted adjacency matrix $A=[a_{ij}]\in\mathbb{R}^{N\times N}$, where
@@
where $w_{ij}>0$ for all $(j,i)\in\mathcal{E}$. For the special case of undirected graphs, $A$ is symmetric and $w_{ij}=w_{ji}$. We define the neighboring agents set $\mathcal{N}_{i}:=\{\,j\in\mathcal{V}:a_{ij}>0\,\}$. The in-degree matrix is denoted by $D=\mathrm{diag}\{d_{i}\}$ and defined via $d_{i}=\sum_{j=1}^{N}a_{ij}$, and the Laplacian is L:=DAL:=D-A.
@@
Under the fixed topology $\mathcal{G}$, the augmented graph $\bar{\mathcal{G}}$ contains a spanning tree with the leader as the root, i.e., the leader is either directly or indirectly connected to every agent. In other words, whenever $(v_{i},v_{0})\notin\bar{\mathcal{E}}$, there exists a sequence of nonzero elements of $A$ of the form ${a}_{ii_{2}},{a}_{i_{2}i_{3}},\cdots,{a}_{i_{l-1}i^{\prime}}$, for some $(i^{\prime},0)\in\bar{\mathcal{E}}$, with the sequence length $l$ being a finite integer. $\square$
To avoid unsafe lack of information in safety critical situations, we impose the following information architecture.
Assumption 3.2
If $\|x^{i}_{1}-x^{j}_{1}\|\leq\phi$ with $\phi\in\mathbb{R}_{>0}$ denoting a position proximity threshold, then $a_{ij}\neq 0$. $\square$
The above assumption ensures that agents within a $\phi$-range are connected and can exchange information.
To accommodate tracking with offsets and ensure coordinated movement while maintaining formation, we denote by $\psi_{p}^{i}\in\mathbb{R}^{d}$ and $\psi_{p}^{0}\in\mathbb{R}^{d}$ desired state offset (formation geometry) from the leader or other agents.
As presented in (Koulong and Pakniyat, 2025, Sections 3.2–3.4) for the distribution of policies, we define the local weighted stability error of each agent $i$ as
@@
Using Theorem 3.1, we formulate the distributed Tube-MPC with eCBF constraints, which can be solved using the continuous-time transcription (multiple shooting with RK4 inside the optimal control problem). At each sampling instant $t_{k}$, we optimize the nominal trajectory $(\bar{x}^{\,i},\bar{u}^{\,i})$ with sampling period $T_{s}$ and horizon length $H$ while enforcing the tightened eCBF inequalities on the nominal states, and apply to the plant the tube feedback. For a given prediction horizon $H\in\mathbb{N}$, we minimize the objective below subject to the dynamics (2), with $k=0,\ldots,H-1$, as follows
@@
where $\Delta\bar{u}^{i}_{k}:=\bar{u}^{i}_{k}-\bar{u}^{i}_{k-1}$ is the control increment that penalizes large control changes between consecutive time steps, $(Q_{r},P_{r},R,R_{\Delta})\succ 0$ are weighting matrices. $\Phi_{ij}^{\text{tight}}$ is given by (32) and $\Phi_{iO}^{\text{tight}}$ is given by (33). The terminal safe set is defined as

### v2

@@ 1 Introduction
To operate successfully in applications like vehicle convoys and aerial swarms, multi-agent systems (MAS) must satisfy strict coordination and safety requirements (Barnes et al., 2007). This challenge becomes more severe for systems with high relative degrees, where chains of integrators between the control input and safety output slow the effective response and increase collision risk in safety-critical settings (Nguyen and Sreenath, 2016; Garg et al., 2024). Model Predictive Control (MPC) provides a powerful framework for constrained control by optimizing over predicted future trajectories (Mayne et al., 2000), while higher-order and exponential control barrier functions (HOCBFs and eCBFs) enable systematic enforcement of safety constraints in systems with complex, high-relative-degree dynamics (Xiao and Belta, 2022, 2019; Tan et al., 2022; Nguyen and Sreenath, 2016; Wang et al., 2021). However, guaranteeing robust safety for MAS under model uncertainty and bounded disturbances remains challenging, particularly when these advanced CBF techniques are deployed in distributed MPC architectures.
@@
This work introduces a robust safety framework for leader–follower formation control in continuous-time, nonlinear multi-agent systems subject to bounded matched disturbances. The central contribution is a method that formally couples robust error feedback with nominal trajectory planning to enforce safety specifications. Our approach complements existing distributed MPC–HOCBF integrations (Wang et al., 2025), whose safety analysis is often confined to nominal, disturbance-free models, and extends beyond other distributed controllers for multi-agent systems that are either reactive and lack the foresight of a prediction horizon (Tan and Dimarogonas, 2022; Mestres et al., 2024) or, when predictive, do not address robustness for high-relative-degree dynamics (Jiang and Guo, 2024). By integrating tube-based robust constructions (Kolathaya and Ames, 2019) with techniques for high-order constraints (Xiao and Belta, 2019; Ames et al., 2017), the proposed framework provides a unified setting that combines predictive planning, robustness to bounded disturbances, and applicability to high-relative-degree multi-agent systems.
Safety-constrained control is often formulated either via barrier Lyapunov functions (BLFs), which enforce constraints through Lyapunov-like functions that diverge at the boundary, e.g., (Koulong and Pakniyat, 2025; Zhao et al., 2025), or via control barrier functions (CBFs), which encode forward invariance as affine-in-control inequalities that integrate naturally with optimization-based controllers such as QP filters and MPC. In this work, we adopt high-relative-degree CBF variants (HOCBF/eCBF) for this tractability and focus on robustness by converting tube-bounded tracking errors into systematic tightenings of the eCBF constraints, so that nominal planning implies safety for the true disturbed system. This perspective is related to robust CBF ideas such as the “tube-CBF” construction of (Schilliger et al., 2021), as well as to single-agent plan–track and safety-filter frameworks including FaSTrack (Herbert et al., 2017), safety filters (Hsu et al., 2024), and funnel libraries (Majumdar and Tedrake, 2017). For multi-agent systems, by contrast, the dominant safety constraints are mutual collision-avoidance conditions of the form $h_{ij}(x_{i},x_{j})\geq 0$, which couple agents’ states and disturbance tubes and cannot be enforced by applying independent copies of a single-agent tube, funnel, or safety filter to each agent. Extending these ideas to multi-agent settings is therefore nontrivial, and robustness must be propagated through joint constraints involving multiple agents’ uncertainty sets; our key contribution is a support-function-based tightening that aggregates per-agent tubes into explicit scalar safety margins for pairwise constraints, enabling distributed MPC feasibility to imply robust safety for continuous-time nonlinear multi-agent systems with high-relative-degree safety outputs.
We propose a control synthesis framework that couples ancillary robust feedback with nominal planning. Each agent uses a local feedback law to confine nominal–true deviations within a precomputed RPI tube, and the tube geometry is then used to systematically tighten the eCBF constraints enforced on the nominal trajectories, so that nominal feasibility implies safety of the true disturbed system. We embed these tightened eCBF constraints in a distributed MPC formulation, where the prediction horizon enables proactive collision/obstacle avoidance and yields a tractable local optimization with end-to-end guarantees of robust safety and recursive feasibility.
The remainder of the paper is organized as follows: Section 2 introduces the dynamics, nominal model, RPI tubes, and support-function-based eCBF tightening; Section 3 presents the DMPC formulation with tightened eCBF constraints under a fixed communication topology; Section 4 provides a numerical validation; and Section 5 concludes with contributions and future directions.
@@ 2.2 Robust Positive Invariant (RPI) Tube Construction:
$\rho_{i}\geq\frac{2\bar{w}_{i}\lambda_{\max}(P_{i})}{\lambda_{\min}(Q_{i})-2L_{f^{i}}\lambda_{\max}(P_{i})},$
where $L_{f^{i}}$ is an upper bound on the Lipchitz constant of $f^{i}$ on a neighborhood of $\bar{x}^{i}$ containing the $\mathcal{Z}_{i}$ variation set. $\square$
@@
it suffices that
@@
We remark that the tube radius $\rho_{i}$ depends on a local Lipschitz constant $L_{f^{i}}$ of $f^{i}$ over a compact region containing the tube $Z_{i}$, and this constant in turn depends on the size of that region. In practice, one can start from a conservative upper bound on $L_{f^{i}}$, compute the resulting tube $Z_{i}$ and radius $\rho_{i}$, and then iteratively refine $L_{f^{i}}$ and $\rho_{i}$ over the compact operating region thus obtained.
@@ 2.3 Tightened Safety and Support Functions
Assumption 2.1
For each pair $(i,j)\in\mathcal{V}\times\mathcal{V}$ and each obstacle $O\in\mathcal{O}$, there exist scalar safety functions $h_{ij}:\mathbb{R}^{n\cdot d}\times\mathbb{R}^{n\cdot d}\to\mathbb{R}$ and $h_{iO}:\mathbb{R}^{n\cdot d}\to\mathbb{R}$ that encode safety via $h_{ij}(x^{i},x^{j})\geq 0$ and $h_{iO}(x^{i})\geq 0$. We assume $h_{ij}$ and $h_{iO}$ are convex and continuously differentiable in their arguments. $\square$
@@
where $F_{x}(x^{i})=\big[\begin{array}[]{ccc}{x}_{2}^{i^{\top}}&\cdots&f^{i}(x^{i},t)^{\top}\end{array}\big]^{\top},$ $F_{u}=[\begin{array}[]{cccc}0&\cdots&0&I\end{array}]^{\top},$ $F_{w}=[\begin{array}[]{cccc}0&\cdots&0&I\end{array}]^{\top}$. Accordingly, the nominal control–affine dynamics of agent $i$ from (2) is written
@@
$\mathcal{C}:=\bigg\{\left(x^{1},\cdots,x^{i},\cdots,x^{N}\right)\in\mathbb{R}^{N\cdot n\cdot d}:\\[-6.0pt] \hfill h_{ij}(x^{i},x^{j})\geq 0,\hskip 20.0pt\forall(i,j)\in\mathcal{V}^{2},\hskip 45.0pt\\[-6.0pt] h_{iO}(x^{i})\geq 0,\hskip 20.0pt\forall(i,O)\in\mathcal{V}\times\mathcal{O}_{i}\bigg\},$ (20)
@@
Given the decomposition $x^{\,i}=\bar{x}^{\,i}+z^{\,i}$, $x^{\,j}=\bar{x}^{\,j}+z^{\,j}$ with $z^{\,i}\in\mathcal{Z}_{i}$, $z^{\,j}\in\mathcal{Z}_{j}$, and $z^{ij}=z^{i}-z^{j}$, we define the inter-agent tightened safety function as the worst-case value (infimum) considering all possible errors in relative tube $\mathcal{Z}_{ij}=\mathcal{Z}_{i}\oplus(-\mathcal{Z}_{j})$ in the form of (illustrated in Figure 1)
@@
Lemma 2.3 (Support Function - Tubes)
The support function of the RPI tube $\mathcal{Z}_{i}:=\{\,z^{\,i}:~(z^{\,i})^{\top}P_{i}z^{\,i}\leq\rho_{i}^{\,2}\,\}$, with $\rho_{i}$ and $P_{i}\succ 0$ satisfying the Lyapunov equation (9) as in Lemma 2.1, is
@@
[…]
Figure 1: Geometric illustration of tube-based tightening for agent–obstacle avoidance in 2D. The nominal state $\bar{x}^{i}(t)$ is surrounded by an RPI tube $Z_{i}$ and $x^{i}(t)$. The control barrier function boundary $h_{iO}$ (black dashed) is tightened by the worst-case decrease of $h_{iO}$ under tube deviations, bounded by $\delta_{iO}(\bar{x}^{i})=\sigma_{Z_{i}}(\nabla_{\bar{x}^{i}}h_{iO})$. Enforcing $h_{iO}(\bar{x}^{i})\geq\delta_{iO}(\bar{x}^{i})$ (equivalently $h_{iO}-\delta_{iO}\geq 0$, orange dashed) guarantees $h_{iO}(x^{i})\geq 0$ for all admissible tube realizations.
@@ 3.1 Proposed Tightened eCBF Functions
We propose the following tightened exponential control barrier functions (tight-eCBF) using the standard eCBF (3)–(3), to impose on the nominal trajectories of agents while ensuring safety of the actual trajectories which are subject to unknown disturbances $w^{i}$ and corresponding errors $z^{i}$ as
@@
where $\kappa_{0},\dots,\kappa_{r-1}>0$ are eCBF gains for a relative-degree-$r$ safety function, and where we use notations $\delta_{ij}(\bar{x}^{i},\bar{x}^{j}):=\sigma_{\mathcal{Z}_{i}}(g_{i})+\sigma_{\mathcal{Z}_{j}}(g_{j})\equiv\rho_{i}\left\|\nabla_{\bar{x}^{i}}h_{ij}(\bar{x}^{i},\bar{x}^{j})\right\|_{P_{i}^{-1}}+\rho_{j}\left\|\nabla_{\bar{x}^{j}}h_{ij}(\bar{x}^{i},\bar{x}^{j})\right\|_{P_{j}^{-1}}$, together with $\delta_{iO}(\bar{x}^{i}):=\sigma_{\mathcal{Z}_{i}}(g_{iO})\equiv{\rho_{i}}\left\|\nabla_{\bar{x}^{i}}h_{iO}(\bar{x}^{i},\bar{x}^{j})\right\|_{P_{i}^{-1}}$.
@@
where $G_{i}^{\max}$ and $G_{j}^{\max}$ are some finite bounds on the gradients of $h_{ij}$ with respect to $\bar{x}^{i}$ and $\bar{x}^{i}$, i.e.,
$\|\nabla_{\bar{x}^{i}}h_{ij}(\bar{x}^{i},\bar{x}^{j})\|\leq G_{i}^{\max},\qquad\|\nabla_{\bar{x}^{j}}h_{ij}(\bar{x}^{i},\bar{x}^{j})\|\leq G_{j}^{\max}$
which exist because $h_{ij}$ is continuously differentiable under Assumption 2.1, and $\bar{x}^{i}$ and $\bar{x}^{i}$ are bounded reference trajectories. Hence
@@ 3.2 Distributed Tube-MPC eCBF Formation Framework
To avoid requiring all-to-all communication, reduce computational load at agent level, and enable distributed control policies, we consider a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E},A)$ with the node set $\mathcal{V}=\{1,\dots,N\}$, edge set $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$, and weighted adjacency matrix $A=[a_{ij}]\in\mathbb{R}^{N\times N}$, where
@@
where $w_{ij}>0$ for all $(j,i)\in\mathcal{E}$. For the special case of undirected graphs, $A$ is symmetric and $w_{ij}=w_{ji}$. We define the neighboring agents set $\mathcal{N}_{i}:=\{\,j\in\mathcal{V}:a_{ij}>0\,\}$.
To enable decentralized communication while avoiding unsafe loss of information in safety-critical situations, we impose the following information architecture.
Assumption 3.1
The communication weights has the property that, if $\|\bar{x}^{i}_{1}-\bar{x}^{j}_{1}\|\leq\phi$ for some position proximity threshold $\phi\in\mathbb{R}_{>0}$, then $a_{ij}\neq 0$. $\square$
The threshold $\phi$, specified as part of the problem setting, is chosen in practice to be no smaller than the safety distance plus the tube-induced safety margins, ensuring that any agents that may approach the safety boundary can exchange the information required for safety enforcement.
The in-degree matrix is denoted by $D=\mathrm{diag}\{d_{i}\}$ and defined via $d_{i}=\sum_{j=1}^{N}a_{ij}$, and the Laplacian is L:=DAL:=D-A.
@@
The augmented graph $\bar{\mathcal{G}}$ contains a spanning tree with the leader as the root, i.e., for every agent $i$ there exists a (directed) path in $\bar{\mathcal{G}}$ from the leader node $v_{0}$ to $v_{i}$. In other words, whenever $(v_{i},v_{0})\notin\bar{\mathcal{E}}$, there exists a sequence of nonzero elements of $A$ of the form ${a}_{ii_{2}},{a}_{i_{2}i_{3}},\cdots,{a}_{i_{l-1}i^{\prime}}$, for some $(i^{\prime},0)\in\bar{\mathcal{E}}$, with the sequence length $l$ being a finite integer. $\square$
To accommodate tracking with offsets and ensure coordinated movement while maintaining formation, we denote by $\psi_{p}^{i}\in\mathbb{R}^{d}$ and $\psi_{p}^{0}\in\mathbb{R}^{d}$ desired state offset (formation geometry) from the leader or other agents. As presented in (Koulong and Pakniyat, 2025) for the distribution of policies, we define the local weighted stability error of each agent $i$ as
@@
Using Theorem 3.1, we pose a DMPC problem with tightened eCBF constraints, solved via continuous-time transcription (multiple shooting with RK4). At each sampling time $t_{k}$, we optimize the nominal sequence $(\bar{x}^{i},\bar{u}^{i})$ over horizon $H$ with step $T_{s}$, enforce the tightened eCBF inequalities on $\bar{x}^{i}$, and apply the tube feedback to the plant. For a given $H\in\mathbb{N}$, we minimize the objective subject to (2) for $k=0,\ldots,H-1$:
@@
where $\Delta\bar{u}^{i}_{k}:=\bar{u}^{i}_{k}-\bar{u}^{i}_{k-1}$, $(Q_{r},P_{r},R,R_{\Delta})\succ 0$ are weighting matrices, and $\Phi_{ij}^{\text{tight}}$, $\Phi_{iO}^{\text{tight}}$ are defined in (32) and (33), respectively. The terminal safe set is defined as